OpenStudy (leahhhmorgannn):

Choose the correct simplification of: x^12z^11 OVER x^2z^4

OpenStudy (leahhhmorgannn):

equation and answer options:

OpenStudy (kendricklamar2014):

Just subtract the exponents:\[12 - 2 = ?\] \[11 - 4 = ?\]

OpenStudy (leahhhmorgannn):

12 - 2 = 10 11 - 4 = 7 then is there anything to do after that?

OpenStudy (leahhhmorgannn):

Is the answer A?

OpenStudy (kendricklamar2014):

Yes the answer is A. Good Job :)

OpenStudy (leahhhmorgannn):

thank you!

OpenStudy (kendricklamar2014):

Your Welcome